# Empirical investigation of multi-source cross-validation in clinical machine learning

### v1

@@ Abstract
In this study, we present a systematic empirical evaluation of standard K-fold cross-validation and leave-source-out cross-validation methods in a multi-source setting. We consider the task of electrocardiogram based cardiovascular disease classification, combining and harmonizing the openly available PhysioNet CinC Challenge 2021 and the Shandong Provincial Hospital datasets for our study.
@@ 1 Introduction
  • We combine and harmonize the openly available PhysioNet/CinC challenge 2021 [11, 12, 13] and the Shandong Provincial Hospital [14] datasets, creating a large multi-source resource for testing the generalization of ECG based diagnostic classifiers for CVDs.
@@
In Section 2 we cover related work on estimating generalization of biomedical classifiers. Section 3 covers the datasets and their harmonization, machine learning methodology, and experimental setup. In Section 4 we present the experimental results, in Section 5 the discussion, and Section 6 concludes the paper.
@@ 2 Related work
Rakotomalala et al. [21] considered the use of CV for multi-source data, using hospital wards as an example. They empirically demonstrated, that if the goal is to estimate generalization performance with respect to new sources, standard K-fold CV produces overoptimistic results. They proposed an alternative approach where CV folds are sampled on the level of data sources, rather than on the level of individual patients. Geras et al. [10] called such approach multiple-source cross-validation, and theoretically showed that it produces an unbiased estimate of the expected prediction performance on data from a new source.
Multiple-source CV has been applied in a number of previous studies. Knight et al. [22] empirically evaluated leave-one-source-out CV for Magnetic Resonance Imaging where data was gathered from different scanners. Han et al. [23] assessed their two models for 12-lead ECG classification, calling multiple-source CV as leave-one-dataset-out CV. Tabe-Bordbar et al. [24] compared the clustering-based CV with the standard CV on the task of gene expression prediction and illustrated how the standard CV produced over-optimistic performance estimates. McWilliams et al. [25] implemented the CV approach with data from two sources for the task of intensive care discharge prediction. They also argued about the challenges of data harmonization when data comes from multiple sources: They chose not to use diagnostic codes and severity of illness, although the features have previously been shown to be useful in predicting adverse events following discharge. Compared to these previous works, our study includes both a greater number of data sources, a greater volume of data (individuals), and more systematic evaluation protocol for assessing the bias, variance and overall mean squared error of the compared cross-validation methods.
@@ 3.1 Data sources
Sources Countries Locations Total patients (n) Total ECGs (n) Included ECGs1 (n) Sampling frequency Length (s) Labeling standard Chapman-Shaoxing and Ningbo China Shaoxing People’s Hospital Ningbo First Hospital 45 152 45 152 43 814 500 10 SNOMED CPSC and CPSC-Extra China 11 unnamed hospitals Unknown 10 330 6 110 500 6-144 SNOMED G12EC USA Emory University Hospital 15 738 10 344 8 892 500 5-10 SNOMED PTB and PTB-XL Germany and other European countries University Clinic Benjamin Franklin Physkalisch Technische Bundesantalt 19 147 22 353 21 348 500, 1000 10-120 SNOMED SPH2 China Shandong Provincial Hospital 24 666 25 770 23 274 500 10-60 AHA
@@
The CinC data, including only recordings with labels of our interest, consist of 80,164 12-lead ECGs from multiple databases. Among these databases, some originate from the same location, which we addressed by integrating them into a unified dataset based on their shared origin. For instance, PhysioNet derived two separate databases from the China Physiological Signal Challenge 2018 (CPSC), the CPSC database and the unused CPSC-Extra database [26, 12]. We treated them as a unified data source, naming it the CPSC and CPSC-Extra dataset. Similar merging was applied to the Physikalisch-Technische Bundesanstalt (PTB) and PTB-XL databases [27, 28] and the Chapman-Shaoxing and Ningbo databases [29, 30]. Consequently, we considered the CinC data as four distinct data sources, the CPSC and CPSC-Extra dataset, the PTB and PTB-XL dataset, the Chapman-Shaoxing and Ningbo dataset and the Georgia 12-lead ECG (G12EC) dataset [31].
@@
The majority of the ECGs in the CinC data are sampled at 500 Hz, but a small number are sampled at 1000 Hz in the PTB and PTB-XL dataset. The duration of the recordings spans from 5 seconds to nearly 2.5 minutes. Most of the recordings include age and gender as demographic information and more detailed lead and recording information. Table 2 describes the age and gender attributes per data source. Whilst the majority of the ECGs include one (44.9 %), two (24.7 %) or three (14.7 %) labels, the total number of labels can vary from 1 to 12, depending on the data source (Figure 1).
@@
The PhysioNet/CinC challenge 2021 used 30 diagnostic labels to evaluate challenge participants’ algorithms. The 30 diagnoses were chosen based on their relative prevalence, clinical significance, and higher likelihood of being identifiable from ECGs [12]. The diagnoses are encoded with approximate Systematized Nomenclature of Medicine Clinical Terms (SNOMED CT) codes.
@@ 3.1.2 Shandong Provincial Hospital Database
The SPH database, released in 2022, was collected at Shandong Provincial Hospital between 2019 and 2020. We treated the SPH data as an independent data source in addition to the four sources that form the PhysioNet/CinC challenge 2021 data. Once the database was filtered to contain only the particular diagnoses chosen for our study, the SPH data contains 23,274 annotated 12-lead ECGs. Of these, 59.7 % of the ECGs are labeled as ”normal ECGs” indicating they do not include any diagnosed abnormalities. As shown in Figure 1, 79.3 % of all the recordings have a single label. 18.5 % have two labels, 1.9 % have three labels, and 0.3 % are assigned either four or five labels.
The ECGs are sampled at 500 Hz with a duration spanning between 10 and 60 seconds. For each ECG, age and gender information is provided (Table 2) as well as the recording date, the number of samples and the patient ID. The portion of the SPH data extracted for our study contains ECGs from 22,334 patients, some of whom have multiple ECGs, albeit the majority of them (21,540) have only a single recording.
Instead of SNOMED CT codes, the SPH data follows the American Heart Association (AHA), for ECG labeling, as recommended by the AHA, the American College of Cardiology, and the Heart Rhythm Society. The labels are provided as primary statements and optional additional modifiers. The primary statements are single labels conveying clinical meaning as is, and the modifiers serve as adjectives to further refine the meaning of the core statement [33].
@@
Source Age Gender1
Mean ($\pm$ SD) Min Max U (n) Ratio (M/F, %)
Chapman-Shaoxing and Ningbo 58.21 ($\pm$ 19.62) 0 89 44 56/43
CPSC and CPSC-Extra 62.69 ($\pm$ 18.67) 1 104 8 57/43
G12EC 60.82 ($\pm$ 15.51) 14 89 65 54/46
PTB and PTB-XL 59.57 ($\pm$ 17.01) 2 95 92 52/48
SPH 49.66 ($\pm$ 15.51) 18 95 0 57/43
Final dataset 57.05 ($\pm$ 18.32) 0 104 209 56/44
  • U = Undefined
  • 1
    Only the Chapman-Shaoxing and Ningbo dataset has 14 ECGs that lack the age value.
@@ 3.1.3 Mapping between SNOMED CT and AHA standards
The CinC data uses the SNOMED CT standard, whereas the SPH data uses the AHA standard for labeling. We created a label mapping to convert AHA codes to SNOMED CT. The mapping was created based on the provided descriptions of the CinC data [11] and the SPH data [14] in collaboration with two cardiologists. The mapping is shown in Table 3.
The mapping has the following nuances. First, in the SPH data, ECGs labeled with a ”normal ECG” label were assigned a value of -1 for easier identification and to prevent their removal. Second, we merged all the prolonged PR interval labels to the 1st degree AV block label, as the 1st degree AV block diagnosis was missing from the SPH data but both were found in the CinC data. Third, as described, the SPH data includes both the primary statements, which can be seen as main diagnostic labels, and the modifiers that refines those statements. For example, atrial fibrillation can be used as a primary statement and it can be further characterized using two modifiers, ”with a rapid ventricular response” or ”with a slow ventricular response”. This results in three distinct diagnostic categories, each containing different ECG recordings. As the SNOMED CT standard does not include such refinements, we merged all the ECGs under atrial fibrillation and its further refinements under ”atrial fibrillation”. Similar merging was applied to ECGs indicating premature atrial contraction, which initially was refined using five additional modifiers.
@@ 3.1.4 Selection of labels
The labels were chosen following specific criteria. The first criterion was that each label must be included in the PhysioNet/CinC challenge 2021 evaluation, i.e., it should be found among the 30 scored diagnoses. These 30 diagnoses were chosen based on their relevance by the PhysioNet challenge team, which included expert cardiologist guidance. The second criterion required that each label had to be present in a minimum of four distinct data sources and occur at least 50 times in each of them. In other words, it was acceptable if a chosen diagnosis was missing or had a low number of instances in one source, as long as it was well-represented in the remaining sources.
Based on the above mentioned criteria, a total of 16 diagnoses along with the sinus rhythm label were selected for ECG classification. Table 3 shows the diagnoses and their respective occurrences in each data source.
@@
Diagnosis SNOMED CT AHA Chapman-Shaoxing and Ningbo CPSC and CPSC-Extra G12EC PTB and PTB-XL SPH Total first degree AV block1 270492004 82 1192 828 769 1012 238 4039 atrial fibrillation 164889003 50 50+346 50+347 1780 1374 570 1529 675 5928 atrial flutter 164890007 51 8060 54 186 74 99 8473 incomplete right bundle branch block 713426002 105 246 86 407 1118 1259 3116 left anterior fascicular block 445118002 101 380 0 180 1626 154 2340 left axis deviation 39732003 121 1545 0 940 5146 138 7769 left bundle branch block 164909002 104 240 274 231 536 84 1365 low QRS voltages 251146004 125 1043 0 374 182 322 1921 premature atrial contraction 284470004 30 30+308 30+310 30+340 30+341 30+349 1312 689 639 398 539 3577 right axis deviation 47665007 120 853 1 83 343 161 1441 right bundle branch block 59118001 106 649 1858 542 0 710 3759 sinus arrhythmia 427393009 23 2550 11 455 772 1553 5341 sinus bradycardia 426177001 22 16559 45 1677 637 2711 21629 sinus rhythm 426783006 -2 8125 922 1752 18172 16858 458292 sinus tachycardia 427084000 21 7255 303 1261 827 725 10371 T wave abnormal 164934002 147 7043 22 2306 2345 2042 13758 T wave inversion 59931005 147+367 2877 5 812 294 176 4164 1 The prolonged PR interval label was merged to the first degree AV block label. 2 Due to the lack of the sinus rhythm label in the SPH data, the labels were imputed.
@@ 3.1.5 Imputation of the sinus rhythm label
Establishing a mapping for the sinus rhythm (SR) label between AHA statements and SNOMED CT codes required a different approach because SR is not explicitly included in AHA statements but is in SNOMED CT codes. The AHA statements clearly differentiate between ”normal” ECGs and those exhibiting abnormalities. Normal ECGs are simply assumed to be in sinus rhythm without any abnormalities. However, in the case of abnormal ECGs, it remains unclear whether the recording is in sinus rhythm. Due to the complexity of diagnoses, we chose to impute the SR label when mapping abnormal ECGs from AHA statements to the SNOMED codes.
@@
The imputation of the SR label to the SPH data included the following steps: i) A one-hot-encoded $n\times 16$ label matrix, where $n$ is the number of the ECGs and the 16 columns correspond to the labels (excluding SR), was constructed both for the CinC and SPH datasets. ii) A logistic regression model ($\lambda$=0.01) was trained on the CinC data to predict the SR label, using the other 16 diagnoses as input features. iii) The logistic regression model was used to impute the SR label to all SPH ECGs labeled as ”abnormal”. The once with ”normal” label were by default assigned the positive SR label. Overall, the SR label was assigned to 16 858 (72.4 %) out of 23 274 ECG recordings in the SPH dataset.
3.2 ECG classifier
The ECG classifier used in this study was originally created for the PhysioNet/CinC challenge 2020 and is implemented in PyTorch [8]. It uses a ResNet architecture with additional squeeze and excitation (SE) blocks.
The classifier consists of a single convolutional layer, followed by eight residual blocks (RBs). Each of these blocks comprises two convolutional layers and the SE block. Both the first convolutional layer and the RB units employ 64 convolutional filters, with the number of filters doubling after every second RB unit. The feature dimension is reduced by half after the max pooling layer and the third, fifth, and seventh RBs. The sigmoid function is applied to each output neuron, each representing one diagnostic label.
Compared to the original ResNet architecture, the SE-ResNet classifier includes four modifications: Firstly, the age and gender information are passed to the network through an additional fully connected layer before being added to the final fully connected layer. Secondly, to enhance the learning of meaningful features, a kernel size of 15 is applied to the first convolutional kernel, while the size of the other kernels is 7. Thirdly, a dropout layer with a dropout rate of 0.2 is placed between the two convolutional kernels in each RB to reduce the risk of overfitting. Finally, a SE block is added into each RB to model the spatial relationships between the ECG channels.
We used binary cross entropy loss and the Adam optimizer with an initial learning rate of 0.003. Each model was trained for 50 epochs with a batch size of 64.

### v2

# Empirical investigation of multi-source cross-validation in clinical ECG classification

@@ Abstract
In this study, we present a systematic empirical evaluation of standard K-fold cross-validation and leave-source-out cross-validation methods in a multi-source setting. We consider the task of electrocardiogram based cardiovascular disease classification, combining and harmonizing the openly available PhysioNet/CinC Challenge 2021 and the Shandong Provincial Hospital datasets for our study.
@@
keywords:
Performance evaluation , Multi-source cross-validation , Electrocardiogram classification , Cardiovascular diseases , Deep learning
journal: Computers in Biology and Medicine
[…]
[utu]organization=Department of Computing, University of Turku,country=Finland
[…]
[leeds]organization=Leeds Institute of Health Sciences, University of Leeds,country=UK
[…]
[leeds_crd]organization=Institute of Cardiovascular and Metabolic Medicine, University of Leeds,country=UK
@@ 1 Introduction
  • 1.
    We combine and harmonize the openly available PhysioNet/CinC challenge 2021 [11, 12, 13] and the Shandong Provincial Hospital [14] datasets, creating a large multi-source resource for testing the generalization of ECG-based diagnostic classifiers for CVDs.
  • 2.
@@
  • 4.
    Finally, we introduce and apply a classification based heuristic for detecting systematic differences between the source distributions.
In this study, we employ four classification algorithms: state-of-the-art deep learning models, including residual networks [8] and transformers [15], alongside logistic regression and XGBoost based baseline methods. To evaluate reliability of different cross-validation approaches, we introduce a comprehensive metric that analyzes mean error (bias), standard deviation, and mean squared error of their estimates. All datasets and code used for these experiments are publicly available at https://github.com/UTU-Health-Research/dl-ecg-classifier. Overall, our work presents the most comprehensive evaluation to date of the source-level cross-validation approach on clinical ECG data.
In Section 2 we cover related work on estimating generalization of biomedical classifiers. Section 3 covers the datasets and their harmonization, machine learning methodologies, and experimental setup. In Section 4 we present the experimental results, in Section 5 the discussion, and Section 6 concludes the paper.
@@ 2 Related work
Rakotomalala et al. [21] considered the use of CV for multi-source data, using hospital wards as an example. They empirically demonstrated, that if the goal is to estimate generalization performance with respect to new sources, standard K-fold CV produces overoptimistic results. They proposed an alternative approach where CV folds are sampled on the level of data sources, rather than on the level of individual patients. Geras et al. [10] called such approach multiple-source cross-validation, and theoretically showed that it produces an unbiased estimate of the expected prediction performance on data from a new source. Multi-source evaluation is closely related to a common scenario in medical data analysis, where data is grouped by subjects, each contributing multiple data points, such as medical scans or biosignal measurements. In these cases, subject-level CV methods, such as leave-subject-out [22, 23], are frequently used to ensure robust model performance across different individuals. Multi-source CV extends this approach to evaluate generalization across entire data sources, such as hospitals, rather than focusing solely on the individual subject level.
Multi-source CV has been applied in a number of previous studies within the medical field. Knight et al. [24] empirically evaluated leave-one-source-out CV for Magnetic Resonance Imaging where data was gathered from different scanners. Han et al. [25] assessed their two models for 12-lead ECG classification, calling multiple-source CV as leave-one-dataset-out CV. Tabe-Bordbar et al. [26] compared the clustering-based CV with the standard CV on the task of gene expression prediction and illustrated how the standard CV produced over-optimistic performance estimates. McWilliams et al. [27] implemented the CV approach with data from two sources for the task of intensive care discharge prediction. Compared to these previous works, our study includes both a greater number of data sources, a greater volume of data, i.e. individuals, and more systematic evaluation protocol for assessing the bias, variance and overall mean squared error of the compared cross-validation methods.
@@ 3.1 Data sources
Sources Countries Locations Total patients (n) Total ECGs (n) Included ECGs (n) Sampling frequency Length (s) Labeling standard Chapman-Shaoxing and Ningbo China Shaoxing People’s Hospital Ningbo First Hospital 45,152 45,152 43,814 500 10 SNOMED CPSC and CPSC-Extra China 11 unnamed hospitals Unknown 10,330 6,110 500 6-144 SNOMED G12EC USA Emory University Hospital 15,738 10,344 8,892 500 5-10 SNOMED PTB and PTB-XL Germany and other European countries University Clinic Benjamin Franklin Physkalisch Technische Bundesantalt 19,147 22,353 21,348 500, 1000 10-120 SNOMED SPH China Shandong Provincial Hospital 24,666 25,770 23,274 500 10-60 AHA
@@
The CinC data, including only recordings with labels of our interest, consist of 80,164 12-lead ECGs from multiple databases. Among these databases, some originate from the same location, which we addressed by integrating them into a unified dataset based on their shared origin. For instance, PhysioNet derived two separate databases from the China Physiological Signal Challenge 2018 (CPSC), the CPSC database and the unused CPSC-Extra database [28, 12]. We treated them as a unified data source, naming it the CPSC and CPSC-Extra source. Similar merging was applied to the Physikalisch-Technische Bundesanstalt (PTB) and PTB-XL databases [29, 30] and the Chapman-Shaoxing and Ningbo databases [31, 32]. Consequently, we considered the CinC data as four distinct data sources, the CPSC and CPSC-Extra source, the PTB and PTB-XL source, the Chapman-Shaoxing and Ningbo source and the Georgia 12-lead ECG (G12EC) source [33].
@@
The majority of the ECGs in the CinC data are sampled at 500 Hz, but a small number are sampled at 1000 Hz in the PTB and PTB-XL source. The duration of the recordings spans from 5 seconds to nearly 2.5 minutes. Most of the recordings include age and sex as demographic information and more detailed lead and recording information. Table 2 describes the age and sex attributes per data source. Whilst the majority of the ECGs include one (44.9 %), two (24.7 %) or three (14.7 %) labels, the total number of labels can vary from 1 to 12, depending on the data source (Figure 1). Table 1 indicates that in the PTB and PTB-XL source, the number of patients is lower than the number of included ECGs in the study, which suggests the presence of multiple recordings for some patients. However, due to the lack of the patient identifiers within the CinC data, we were unable to examine the potential for duplicate recordings per patient.
@@ 3.1.2 Shandong Provincial Hospital Database
The SPH database, released in 2022, was collected at Shandong Provincial Hospital between 2019 and 2020. We treated it as an independent data source in addition to the four sources that form the PhysioNet/CinC challenge 2021 data. Once the data source was filtered to contain only the particular diagnoses chosen for our study, the SPH source contains 23,274 annotated 12-lead ECGs. Of these, 59.7 % of the ECGs are labeled as ”normal ECGs” indicating they do not include any diagnosed abnormalities. As shown in Figure 1, 79.3 % of all the recordings have a single label. 18.5 % have two labels, 1.9 % have three labels, and 0.3 % are assigned either four or five labels.
The ECGs are sampled at 500 Hz with a duration spanning between 10 and 60 seconds. For each ECG, age and sex information is provided (Table 2) as well as the recording date, the number of samples, and the patient ID. The portion of the SPH source extracted for our study contains ECGs from 22,334 patients, some of whom have multiple ECGs, albeit the majority of them (21,540) have only a single recording. We compared the shared metadata of the ECGs per patient and noticed that multiple ECGs shared identical metadata, including diagnoses, patient ID, age, sex, the number of samples, and the date the ECG was recorded. In total, 144 ECGs had duplicates based on the metadata alone, and of these, 132 (0.13 % of the final dataset) had identical ECG recording data.
Instead of SNOMED CT codes, the SPH source follows the American Heart Association (AHA), for ECG labeling, as recommended by the AHA, the American College of Cardiology, and the Heart Rhythm Society. The labels are provided as primary statements and optional additional modifiers. The primary statements are single labels conveying clinical meaning as is, and the modifiers serve as adjectives to further refine the meaning of the core statement [35].
@@
Source Age Sex Mean ($\pm$ SD) Min Max U (n) Ratio % (Male/Female) Chapman-Shaoxing and Ningbo 58.21 ($\pm$ 19.62) 0 89 44 56/43 CPSC and CPSC-Extra 62.69 ($\pm$ 18.67) 1 104 8 57/43 G12EC 60.82 ($\pm$ 15.51) 14 89 65 54/46 PTB and PTB-XL 59.57 ($\pm$ 17.01) 2 95 92 52/48 SPH 49.66 ($\pm$ 15.51) 18 95 0 57/43 Final dataset 57.05 ($\pm$ 18.32) 0 104 209 56/44 U = Undefined |||| The Chapman-Shaoxing and Ningbo source has 14 ECGs that lack the sex value.
@@ 3.1.3 Mapping between SNOMED CT and AHA standards
The CinC data uses the SNOMED CT standard, whereas the SPH source uses the AHA standard for labeling. We created a label mapping to convert AHA codes to SNOMED CT. The mapping was created based on the provided descriptions of the CinC data [11] and the SPH source [14] in collaboration with two cardiologists. The mapping is shown in Table 3.
The mapping has the following nuances. First, in the SPH source, ECGs labeled with a ”normal ECG” label were assigned a value of -1 for easier identification and to prevent their removal. Second, we merged all the prolonged PR interval labels to the first degree atrioventricular (AV) block label, as the first degree AV block diagnosis was missing from the SPH source but both were found in the CinC data. Third, as described, the SPH source includes both the primary statements, which can be seen as main diagnostic labels, and the modifiers that refines those statements. For example, atrial fibrillation can be used as a primary statement and it can be further characterized using two modifiers, ”with a rapid ventricular response” or ”with a slow ventricular response”. This results in three distinct diagnostic categories, each containing different ECG recordings. As the SNOMED CT standard does not include such refinements, we merged all the ECGs under atrial fibrillation and its further refinements under ”atrial fibrillation”. Similar merging was applied to ECGs indicating premature atrial contraction, which initially was refined using five additional modifiers.
@@ 3.1.4 Selection of labels
The labels were chosen following specific criteria. The first criterion was that each label must be included in the PhysioNet/CinC challenge 2021 evaluation, i.e. it should be found among the 30 scored diagnoses. These 30 diagnoses were chosen based on their relevance by the PhysioNet challenge team, which included expert cardiologist guidance. The second criterion required that each label had to be present in a minimum of four distinct data sources and occur at least 50 times in each of them. In other words, it was acceptable if a chosen diagnosis was missing or had a low number of instances in one source, as long as it was well-represented in the remaining sources.
Based on the above mentioned criteria, a total of 16 diagnoses along with the sinus rhythm label were selected for ECG classification. Table 3 shows the diagnoses and their respective occurrences in each data source. The data sources are highly imbalanced, with significant differences in label distributions across each source. A key objective of this study is to simulate real-world conditions where data is aggregated from multiple hospitals and to examine how these inherent disparities impact model generalization and performance evaluation.
@@
Diagnosis SNOMED CT AHA Chapman-Shaoxing and Ningbo CPSC and CPSC-Extra G12EC PTB and PTB-XL SPH Total first degree AV block1 270492004 82 1192 828 769 1012 238 4039 atrial fibrillation 164889003 50 50+346 50+347 1780 1374 570 1529 675 5928 atrial flutter 164890007 51 8060 54 186 74 99 8473 incomplete right bundle branch block 713426002 105 246 86 407 1118 1259 3116 left anterior fascicular block 445118002 101 380 0 180 1626 154 2340 left axis deviation 39732003 121 1545 0 940 5146 138 7769 left bundle branch block 164909002 104 240 274 231 536 84 1365 low QRS voltages 251146004 125 1043 0 374 182 322 1921 premature atrial contraction 284470004 30 30+308 30+310 30+340 30+341 30+349 1312 689 639 398 539 3577 right axis deviation 47665007 120 853 1 83 343 161 1441 right bundle branch block 59118001 106 649 1858 542 0 710 3759 sinus arrhythmia 427393009 23 2550 11 455 772 1553 5341 sinus bradycardia 426177001 22 16559 45 1677 637 2711 21629 sinus rhythm 426783006 -2 8125 922 1752 18172 168582 45829 sinus tachycardia 427084000 21 7255 303 1261 827 725 10371 T wave abnormal 164934002 147 7043 22 2306 2345 2042 13758 T wave inversion 59931005 147+367 2877 5 812 294 176 4164 1 The prolonged PR interval label was merged to the first degree AV block label. 2 Due to the absence of the sinus rhythm label in the SPH data, the labels were imputed.
@@ 3.1.5 Imputation of the sinus rhythm label
Establishing a mapping for the sinus rhythm (SR) label between AHA statements and SNOMED CT codes required a different approach because SR is not explicitly included in AHA statements but is in SNOMED CT codes. The AHA statements clearly differentiate between ”normal” ECGs and those exhibiting abnormalities. Normal ECGs are simply assumed to be in sinus rhythm without any abnormalities. However, in the case of abnormal ECGs, it remains unclear whether the recording is in sinus rhythm. Due to the complexity of diagnoses, we chose to impute the SR label when mapping abnormal ECGs from the AHA statements to the SNOMED CT codes.
@@
The imputation of the SR label to the SPH source included the following steps: 1) A one-hot-encoded $n\times 16$ label matrix, where $n$ is the number of the ECGs and the 16 columns correspond to the labels (excluding SR), was constructed both for the CinC and SPH sources. 2) A logistic regression model ($\lambda$=0.01) was trained on the CinC data to predict the SR label, using the other 16 diagnoses as input features. 3) The logistic regression model was used to impute the SR label to ECGs of the SPH source labeled as ”abnormal”. The once with ”normal” label were by default assigned the positive SR label. Overall, the SR label was assigned to 16,858 (72.4 %) out of 23,274 ECG recordings in the SPH source.
3.2 Predictive models
For ECG classification, we used two deep neural network architectures and two traditional machine learning methods: a residual neural network, a transformer neural network, logistic regression (LR), and XGBoost. The neural network architectures include the top two performers from the PhysioNet/CinC Challenge 2020: the residual neural network is based on the second-place entry [8] and the transformer neural network on the winning entry [36].